Please try to help me on this question:
If the measure of an angle is (7x-13)degree and the measure of its supplement is (5x-11)degree, find x.
: Please try to help me on this question:
If the measure of an angle is (7x-13)degree and the measure of its supplement is (5x-11)degree, find x.

Answer by jim_thompson5910(9165)

Since the two angles are supplementary, this means that the two angles add to 180 degrees. So this means that (7x-13)+(5x-11)=180


7x-13+5x-11=180 Start with the given equation.


12x-24=180 Combine like terms on the left side.


12x=180+24 Add 24 to both sides.


12x=204 Combine like terms on the right side.


x=(204)/(12) Divide both sides by 12 to isolate x.


x=17 Reduce.


----------------------------------------------------------------------

Answer:

So the answer is x=17